OpenStudy (anonymous):

Georgia built a model of the Empire State Building that is proportional to the actual building. The actual building is 187 feet wide and 1,454 feet tall. Her model is 10 feet tall. Approximately how wide is Georgia's model?

OpenStudy (anonymous):

if the ratio is the same 1454/10=187/x x =1/286 feet wide